OpenStudy (anonymous):

Which statement best describes the graph of -x^3 - x^2 + 4x + 4? It starts down on the left and goes up on the right and intersects the axis at x = -2, 2, and 4. It starts down on the left and goes up on the right and intersects the axis at x = -2, -1, and 2. It starts up on the left and goes down on the right and intersects the x-axis at x = -2, -1, and 2. It starts up on the left and goes down on the right and intersects the axis at x = -2, 2, and 4.

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (campbell_st):

well simply substitute the x values to find which ones give y = 0 remember that if x = -2 is a zero then (x - 2) is a factor so looking at the 1st choice x = -2 (x +2) x = 2 so (x - 2) and x = 4 so (x - 4) then the product of 2, -2 and -4 needs to be 4 from the equation these multiply to 16 an alternate method is to graph the equation using software... https://www.desmos.com/calculator when you type it in start with y = ... your equation... match what you see to the answers